FEL Upgrade Project Weekly Brief

April 8-12, 2002



A possible explanation was uncovered this week as to why we had difficulty using the multi-slit upstream of the linac in the IR Demo machine. Without a bellows it was most likely not well enough aligned to get the beam through it.

Preliminary testing started this week on the 300 kW HVPS’s for the Injector klystrons.


Preparations continue for the upcoming semiannual project review on May 1-3.

WBS 3 (Beam Physics):

Still slogging through field quality related issues for dipoles and quads. Documentation of dipoles is in progress and discussions of quad powering are underway (EPICS asynchronicity makes this entertaining). Thanks to Jay Benesch for reopening this can of worms and Mike Tiefenback for sharing his considerable experience with CEBAF (sort of a "by his stripes we are healed" kind of thing).

Review prep spooling up, encouraging more thought about commissioning. Discussion of space (or lack thereof) for the multi-slit diagnostic immediately upstream of the linac led Benson to make the very amusing observation that the device, as presently installed, is not on a bellows and thus may not be very well aligned. This could go a long way towards explaining why its hard to fit beam through it. Were Jock still here, he'd certainly award Steve this week's Blue Star award for that call!!!

WBS 4 (Injector):

Work continued on the support tube and ball cathode. The test set up for the W&M implant system is ready to be installed. Adjustments to the shielded bellows/BPM design are being done. The gun stand and the extractor cart have been delivered.

The 5" diameter ceramic cylinders for the Conditioning Resistor were received this week. A holding fixture was designed and given to the Machine Shop to build. The design of the piece parts for the combined Conditioning and Running Resistor mechanism continues. The old Gun HVPS Tank is to be removed from the vault this weekend.



WBS 6 (RF):

Quarter HVPS - Applied 208 VAC to the HVPS for position 3 and started testing the control circuits. Several errors were found and corrected. These changes are being incorporated in the other HVPS. The routing and placement of the HV wiring is nearly complete for unit 3.

Quarter Klystrons - Parts are still missing for the LCW piping on the klystron carts, but is due later today. The Machine Shop is ready to start the work.

WBS 8 (Instrumentation):

The fabrication of the individual beam viewer & camera interface boxes is nearly complete. The interface boxes provide the interface to the control system, the Machine Protection System (MPS) and power for the cameras and lamp rings. Seven (7) out of 36 new 3" shielded viewers (for both IR & UV machines) remain to be wired. During the construction of the last FEL this work was done post installation in the accelerator enclosure, by doing it in the lab there is ample time for the students to do the wiring, and same time during commissioning. Each of these are fully tested and signed off then sent on to the vacuum group for girder installation. The few viewers that remained installed will be refitted with the new interface as time allows. The new software screens for viewer and video control are nearly complete. These will be tested beginning next week.

The design is complete for the compact 1.75" injector BPM. The new design has a slot length of ~3" which should allow easy installation with bolts (not studs) and enable the use of a standard bellows upstream. The channel assignments for Switched Electrode Electronics (SEE) and 4 channel electronics have also been made. The software changes will be done in a few weeks.

The design phase of the Beam Loss Monitors (BLM) is nearing completion, the system print block drawings have been started, as well as the wiring diagram for the BLM's. The connectors, voltage dividers and photomultipliers for the additional BLMs have been ordered. The rip out of the damaged Vacuum System HV cables in the vault is near completion, also other unidentifiable cable in the vault was ripped out. The new high voltage cable and connectors have been ordered.

The order for the server upgrade is being written and passed around for signoff. The existing HP servers will be replaced by a SUN and a HP. The Sun will act as a file server and migrate along to replace the HP completely as the rest of the accelerator site computers are changed over. 5 additional (refurbished) workstations were order to upgrade the control room consoles. Additionally 4 flat screen monitors were ordered for the control room to clear more desk space and lessen the magnetic interference from the color monitors (TVs). Other VME control hardware is advancing through production. A prototype of the VME Interface Module (F0117) is being assembled in-house, a prototype of the MPS fiber driver/input module (F0118) is being manufactured, production quantities (24) of the Altera programmable logic board are being fabricated.







WBS 9 (Transport):


Optical Chicane Dipoles (GW)

• At Magnet Enterprises International (MEI) in Oakland CA. The fourth coil looks very good. The fifth is in potting.

• Return legs remain in final machining. Grinding of 4 top and 4 bottom slabs is complete as a batch. The machining subcontractor promises that this batch will be machined "by the end of the month". Depending on how smooth the machining goes there is a possibility the trial assembly of the first article core and parts may happen by the Review on May 1. We still have to glue the pole face shims in place and that will not be complete.

Injector Dipoles (DU/DV)

• Master Machine continues rough machining the slabs for the yokes. Covers are complete. Field clamp material is being cut

• WANG NMR wound all the GU and GV Coils. They are being ground wrapped in fiberglass tape. Epoxy is due on Monday.

Arc Dipoles (GY, GX, GQ)

• The steel for GQ-GX magnet cores was shipped to Process Equipment Co.

• The copper for the GQ-GX coils was received by Wang NMR.

• AES continues to magnetically model the GQ-GX dipoles at lower energy with shorter pole tips to match David Douglas’ use matrix.

• Wang NMR is ready to wind the GY coil as soon as they wrap the insulation around the conductor. They are making the potting mold for the GY coils. They were delayed in starting impregnation of the GG coils by late delivery of epoxy.

• Bosma Machine has the GY core pieces in final planning of the pole surface and the return legs in finish grinding. All ancillary parts are complete


3 inch quad (QX)

• We decided to use the Bang-bang Method of running through hysteresis loops advocated by Mike Tiefenback, magnet integrator for CEBAF. Our production measurements have not used this hysteresis cycle and we are testing several magnets to determine if there is a constant offset to the data that can be applied to our existing readings.

• The power supply being used to test the quads was observed to have an unacceptable ripple. At week’s end, due to these questions, we have developed a circuit to smooth the ripple. We are also making sure the power supply has no regulation problems. We then plan to continue the measurements.

Trim Quad (QT)

• Milhous Control of Virginia reports that potting a first article continues to be delayed because the replacement batch of epoxy is still in transit.

Sextupole (SF)

• DULY Research reports excellent characteristics of its simplified magnetic model. Mechanical design continued with the conclusion that the coils have to be wound two-in-hand in order to have two cooling circuits. The complete concept of how coil hook up would be carried out was then generated.

Octupole (OT)

• Work on this magnet is on furlough until the sextupole is designed.

Beam Line and Vacuum

• Design of the ARC Chambers at AES continues. A full set of drawings, ready for final check is planned for next week.

• The final review of the assembly and detail drawings of the girders for the six quadrupole telescope in front of the first arc was completed.

• The Region in front of the first cryomodules including the injection line went through its first review.

WBS 11 (Optics):

Bids are being solicited for the new vacuum valves for the optical cavity chambers and optical transport system. We anticipate receiving the bids on the optical cavity internals today (4/12) and making the award. Detailing of the lead screw holding assembly is nearly complete and should be completed next week. As the designer for the vacuum vessels is unavailable, that effort was reassigned. We held a meeting to bring the new designer up to speed and are making good progress. Modeling of the 50kW laser beam dump has been extended to the mesh design. We plan to do most of the fabrication in-house, so drawings are being reviewed by our shop now. Design of the chamber and optics for the THz light pickoff continue. We are evaluating using an ultraviewer form the IR Demo to actuate the pickoff mirror.

We moved the pump laser for the ultrafast laser system into UL 1, where it will be initially checked out, once the LSOP is modified and approved. We have discovered that the lens that couples the drive laser output into the single-mode fiber is creating diffraction rings because of the diamond tooling process. This is unacceptable for our pulse compressor, so we are looking at other vendors. Work continues on the OCMMS piezo transducer controls. The mechanical mount has resonance at ~ 50, 110, and 150 Hz, and so the feedback circuit frequency response has to be lowered to avoid them. We are adding a transducer to the mirror mount to mimic low frequency building vibrations, and then watching how well the circuit counteracts them. The OCMMS mirrors are being procured. We are soliciting a new round of estimates for the other cavity optics and the transport optics.

Other activities:

We provided some additional assistance to the Physics Div. with their Compton polarimeter rebuild. We held a kick-off meeting on the DIRUP-funded (through the College of William and Mary) work we will do to support these valuable additions to the facility. I did additional work to prepare for the May 3rd review.

Work continues on the design of the cryocooled mirrors. Finite-element analyses indicates that we need to make the mirrors thicker than originally planned in order to keep the temperature rise low and thus keep the mirror in the temperature region where the thermo-mechanical properties are optimum (~ 50 K).